OpenStudy (tiffany_rhodes):

I know that Euler's number e is defined as the limit of (1+1/x)^x as x approaches infinity.

OpenStudy (tiffany_rhodes):

I would like to know how to solve this limit (without the use of calculus). I've done part of it.

OpenStudy (tiffany_rhodes):

\[\lim_{x \rightarrow \infty} (1+\frac{ 4.8 }{ x })^x\]

OpenStudy (tiffany_rhodes):

I know the answer is e^(4.8)

OpenStudy (tiffany_rhodes):

This is my work so far: \[\ln((1+\frac{ 4.8 }{ x })^x) = x \times \ln(1+\frac{ 4.8 }{ x})\] You could then use one of the limit laws that says that the limit of the products is the product of the limits. \[\lim_{x \rightarrow \infty} x \times \lim_{x \rightarrow \infty} \ln(1+\frac{ 4.8 }{ x })\] The the first limit is infinity multiplied by the second is ln(1) =0. I took the limit inside because of continuity.

OpenStudy (tiffany_rhodes):

Then you would see \[\infty \times 0\] which I feel would qualify as an indeterminate form. That's where I don't know how to proceed.

OpenStudy (perl):

we can use Lhopitals rule now

OpenStudy (perl):

but first you need to make an indeterminate form of \( \large \frac {\infty } {\infty} \) or \( \large \frac {0 } {0} \)

OpenStudy (tiffany_rhodes):

Hmm, yeah that's where I got confused. I'm not quite sure how to change it into that form, unless you could divide one limit by the other ?

OpenStudy (tiffany_rhodes):

In which case that would be infinity divided by 0 or vise versa, which isn't what we want.

OpenStudy (perl):

$$\Large {\lim_{x \to \infty }\ln((1+\frac{ 4.8 }{ x })^x) \\ \lim_{x \to \infty }= x \cdot \ln(1+\frac{ 4.8 }{ x}) \\ \lim_{x \to \infty }= \frac{ \ln(1+\frac{ 4.8 }{ x})}{\frac1x} }$$

OpenStudy (perl):

because x = 1 / (1/ x )

OpenStudy (perl):

$$ \Large { x= \frac{1}{\frac1x }} $$

OpenStudy (tiffany_rhodes):

and the limit would be of the indeterminate form infinity over infinity? So we could then use l'hospitals rule and take the derivative of the top and bottom?

OpenStudy (perl):

it would be indeterminate form of 0/0

OpenStudy (perl):

$$ \Large {\lim_{x \to \infty }\ln((1+\frac{ 4.8 }{ x })^x) \\ \lim_{x \to \infty }= x \cdot \ln(1+\frac{ 4.8 }{ x}) \\ \lim_{x \to \infty }= \frac{ \ln(1+\frac{ 4.8 }{ x})}{\frac1x}\\ ~\\ =\frac{ \ln(1+\frac{ 4.8 }{ \infty})}{\frac{1}{\infty}} = \frac{\ln(1 + 0)}{0}= \frac{\ln(1)}{0}= \frac{0}{0} } $$

OpenStudy (tiffany_rhodes):

Alright, I'm following so far.

OpenStudy (perl):

and therefore we can use Lhopitals rule :)

OpenStudy (tiffany_rhodes):

alright, thanks a lot :)

OpenStudy (perl):

did you do Lhopitals rule?

OpenStudy (tiffany_rhodes):

Yes, sorry got sidetracked lol. I believe I got ln * 4.8

OpenStudy (perl):

$$ \Large {\lim_{x \to \infty }\ln((1+\frac{ 4.8 }{ x })^x) \\ \lim_{x \to \infty }= x \cdot \ln(1+\frac{ 4.8 }{ x}) \\ \lim_{x \to \infty }= \frac{ \ln(1+\frac{ 4.8 }{ x})}{\frac1x} \\ \huge \Rightarrow^{Lhopitals}\\ \LARGE \lim_{x \to \infty } \frac{ \frac{1 }{(1+\frac{ 4.8 }{ x}) } \cdot \frac{-4.8}{x^2} }{-\frac{1}{x^2}} \\ = \lim_{x \to \infty } \frac{1 }{(1+\frac{ 4.8 }{ x}) } \cdot \frac{-4.8}{x^2} \cdot\frac{-x^2}{1} \\ = \lim_{x \to \infty } \frac{1 }{(1+\frac{ 4.8 }{ x}) } \cdot \frac{-4.8}{\cancel{x^2}} \cdot\frac{-\cancel{x^2}}{1} } $$

OpenStudy (tiffany_rhodes):

Oh yeah, chain rule.

OpenStudy (tiffany_rhodes):

Thanks so much @perl I appreciate how detailed your explanation was. It was very helpful.

OpenStudy (perl):

$$ \Large \rm { Suppose~ that~ the~ limit ~ exists,~ call~ it ~ L \\ L = \lim_{x \rightarrow \infty} (1+\frac{ 4.8 }{ x })^x \\~\\ Then ~ it~ follows~ \\ \ln(L) = \ln \left( \lim_{x \rightarrow \infty} (1+\frac{ 4.8 }{ x })^x \right) \\=\lim_{x \to \infty }\ln(1+\frac{ 4.8 }{ x })^x \\ =\lim_{x \to \infty } x \cdot \ln(1+\frac{ 4.8 }{ x}) \\ =\lim_{x \to \infty } \frac{ \ln(1+\frac{ 4.8 }{ x})}{\frac1x} \\ \huge \Rightarrow^{Lhopitals}\\ \LARGE =\lim_{x \to \infty } \frac{ \frac{1 }{(1+\frac{ 4.8 }{ x}) } \cdot \frac{-4.8}{x^2} }{-\frac{1}{x^2}} \\ = \lim_{x \to \infty } \frac{1 }{(1+\frac{ 4.8 }{ x}) } \cdot \frac{-4.8}{x^2} \cdot\frac{-x^2}{1} \\ = \lim_{x \to \infty } \frac{1 }{(1+\frac{ 4.8 }{ x}) } \cdot \frac{-4.8}{\cancel{x^2}} \cdot\frac{-\cancel{x^2}}{1} \\ = \lim_{x \to \infty } \frac{4.8 }{(1+\frac{ 4.8 }{ x}) } \\~\\ = \frac{4.8}{(1+\frac{ 4.8 }{ \infty }) } = \frac{4.8}{(1+0 )}= \frac{4.8}{1} =4.8 \\~\\ \therefore \\~\\ \\ \ln(L) = 4.8 ~ \iff L = e^{4.8} } $$

OpenStudy (perl):

your welcome :)

OpenStudy (tiffany_rhodes):

:)